OpenStudy (anonymous):

Write the point-slope form of the equation of the line passing through the point (-12,-5) with the slope of 3. (1 point)

OpenStudy (anonymous):

y+5=3(x+12)

OpenStudy (anonymous):

Thank you so much☝️

OpenStudy (anonymous):

See \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

Do you mind if I ask you a couple more questions?